If a monopolist faces a demand curve given by the equation P=200-5Q, and their MC=ATC=20, then the consumer surplus resulting from the monopoly price will be...

a. $660

b. $880

c. $990

d. $770
Answer #1

At eqm in MONOPOLY

MR = MC

200-10Q = 20

Q* = 180/10 = 18

P* = 200-5*18 = 110

.

CS = .5*(200-110)*18

= .5*90*18

= $ 810

so correct option not provided